# Intrinsic training dynamics of deep neural networks

### v1

@@ Abstract
A fundamental challenge in the theory of deep learning is to understand whether gradient-based training in high-dimensional parameter spaces can be captured by simpler, lower-dimensional structures, leading to so-called implicit bias. As a stepping stone, we study when a gradient flow on a high-dimensional variable $\theta$ implies an intrinsic gradient flow on a lower-dimensional variable $z=\phi(\theta)$, for an architecture-related function $\phi$. We express a so-called intrinsic dynamic property and show how it is related to the study of conservation laws associated with the factorization $\phi$. This leads to a simple criterion based on the inclusion of kernels of linear maps which yields a necessary condition for this property to hold. We then apply our theory to general ReLU networks of arbitrary depth and show that, for any initialization, it is possible to rewrite the flow as an intrinsic dynamic in lower dimension that depends only on $z$ and the initialization, when $\phi$ is the so-called path-lifting. In the case of linear networks with $\phi$ the product of weight matrices, so-called balanced initializations are also known to enable such a dimensionality reduction; we generalize this result to a broader class of relaxed balanced initializations, showing that, in certain configurations, these are the only initializations that ensure the intrinsic dynamic property. Finally, for the linear neural ODE associated to the limit of infinitely deep linear networks, with relaxed balanced initialization, we explicitly express the corresponding intrinsic dynamics.
@@ 1 Introduction
A central question in deep learning theory is how the complexity of high-dimensional gradient-based training can give rise to simpler, lower-dimensional dynamics. In this work, we explore when the gradient flow on high-dimensional parameters $\theta$ naturally induces a gradient flow on a lower-dimensional variable $z=\phi(\theta)$, where $\phi$ captures structural aspects of the model.
Intrinsic lower dimensional flow
@@
The study of high-dimensional optimization flows arising in the training of neural networks often benefits from the identification of lower-dimensional intrinsic dynamics. Specifically, it is of considerable interest to rewrite a high-dimensional parameter flow $\theta(t)$ in terms of an lower-dimensional representation $z(t)=\phi(\theta(t))$, using a suitable architecture-related reparametrization $\phi$. When the reduced flow on $z(t)$ admits a mirror flow representation, one gains access to powerful analytical tools rooted in convex optimization theory, allowing explicit characterization of the implicit bias induced by gradient-based optimization algorithms. In particular, prior research has successfully leveraged mirror flow formulations to rigorously demonstrate implicit regularization effects, such as sparsity in scalar linear neural networks and two-layer networks with a single neuron (Gunasekar et al. (2018)), as well as maximum-margin classification for logistic regression problems in separable data scenarios (Soudry et al. (2018); Chizat & Bach (2020)). Recent work by Li et al. (2022) identifies sufficient conditions under which the reparametrization $\phi$ makes the flow a mirror flow, requiring that the parametrization be commuting. However, this commuting condition is rarely satisfied in practical scenarios. A weaker condition, introduced by Marcotte et al. (2023), demands merely that the parametrization be involutive, thus ensuring that the flow on $z(t)$ can still be expressed as a Riemannian gradient flow (but not necessarly a mirror flow anymore) dependent solely on $z$ and the initial parameters $\theta(0)$. Marcotte et al. (2023) have shown that this weaker condition applies specifically to the parametrization used in two-layer ReLU networks (Stock & Gribonval (2022)).
Conservation laws
@@
The functions conserved during the dynamics play a crucial role in rewriting flows in lower-dimensional representations. Indeed, when a trajectory $\theta(t)$ is constrained to remain within level sets $\{\theta:\mathbf{h}(\theta)=\mathbf{h}(\theta_{0})\}$ where $\mathbf{h}$ is a (vector-valued) conserved function, the dynamics are effectively restricted to a manifold of lower dimension. A particularly important class of conserved functions along these trajectories is given by the conservation laws associated with a certain architecture-dependent parametrization $\phi$, a concept introduced in Marcotte et al. (2023). These laws depend exclusively on $\phi$, and notably, in the context of neural network training dynamics, they represent quantities preserved across trajectories irrespective of the initial conditions or the training data-set. In the specific case of linear and ReLU neural networks, these conservation laws correspond exactly to previously known “canonical” conserved functions identified in Du et al. (2018), as demonstrated by Marcotte et al. (2023). Furthermore, Marcotte et al. (2023) establish that if the parametrization $\phi$ is involutive, there exist sufficiently many scalar conservation laws to fully rewrite the original trajectory $\theta(t)$ in terms of $\phi(\theta(t))$ and the initial conditions alone. In the linear network case, when so-called balanced conditions (a notion introduced in Arora et al. (2019)) are satisfied (i.e., when the initialization sets all canonical conservation laws (Chitour et al. (2018)) to zero), it becomes possible to rewrite the flow in terms of $z=\phi(\theta)$, where $\phi$ corresponds to the product of weight matrices, as a Riemannian metric that only depend on $z$ and $\mathbf{h}(\theta_{0})=0$ (Arora et al. (2018); Bah et al. (2022)). Moreover, Achour et al. (2025) extended this result to linear convolutional networks in the case of a mean squared loss, but this time for arbitrary initializations, with the Riemannian metric depending on the initialization. For linear networks and in the particular case when the loss function is the square loss, Bah et al. (2022) show that the trajectory evolves on the manifold of matrices having some fixed rank under balanced condition. Still in the square-loss setting, and in the case of two-layer linear networks, Tarmoun et al. (2021); Braun et al. (2022); Dominé et al. (2025) exploit the conservation laws to obtain an exact closed-form expression for $z(t)$ under specific configurations, whereas Varre et al. (2023) uses the same laws to analyse an implicit bias of this dynamic.
Our main contributions.
@@
We first define the notion of intrinsic dynamic property (Definition˜2.6), then the notion of intrinsic metric property (Definition˜2.10) and finally the one of intrinsic recoverability property (Definition˜3.1), and we show the implications (Lemma˜2.11 and Lemma˜3.2):
@@
We then provide a simple criterion that characterizes the intrinsic recoverability property (Theorem˜3.3), and show (Proposition˜3.7) that this criterion is quasi equivalent to the Frobenius property (Definition˜3.6). We prove that the reparametrization for general ReLU networks of arbitrary depth satisfies this property (Theorem˜3.8), establishing that any initialization of a general ReLU network satisfies the intrinsic recoverability property (Corollary˜3.9), as illustrated by a characterization of the intrinsic dynamic of a 3-layer neural network (Proposition˜3.10). Next, by establishing a necessary condition for the intrinsic metric property to hold based on the study of kernels of linear mappings Theorem˜2.14, we show that the intrinsic metric property fails to hold for the natural reparametrizations corresponding of 2-layer linear networks (resp. of attention layers), unless the initialization satisfies the relaxed balanced condition introduced in Definition˜4.1 (Theorem˜4.4). We then show that relaxed balanced initializations do satisfy the intrinsic metric property, not only in 2-layer networks (Theorem˜4.3) but also in linear networks of arbitrary depth (Theorem˜4.6), and we characterize the resulting intrinsic dynamic. Finally, we extend our analysis to the infinite-depth limit of linear networks. We show that a set of functions is conserved along the trajectory (Proposition˜4.7), and, in contrast to the case $L>2$-layer, we derive a closed-form expression for the metric in the case of relaxed balanced initializations (Theorem˜4.8).
@@ 2 Dynamics of over-parameterized models
In most machine learning models, overparameterization occurs due to inherent symmetries (such as rescaling) within the parameter space $\theta\in\mathbb{R}^{D}$. In practice, this redundancy can be factored out through a function $\phi$ (often called a lifting (Candès et al., 2013; Gonon et al., 2023)) that captures these symmetries. Although the resulting lifted variable $z=\phi(\theta)\in\mathbb{R}^{d}$ often lives in higher dimension $d\gg D$, it also belongs to a lower dimensional manifold $\mathcal{Z}$ of dimension $d^{\prime}<D$, and provides a representation of the essential structure of the model.
We consider parameters $\theta(t)\in\mathbb{R}^{D}$ that evolve according to the gradient flow dynamic:
$\dot{\theta}(t)=-\nabla\ell(\theta(t)),$ (1)
to minimize the function $\ell$. In machine learning, $\ell(\theta)$ is typically defined as the empirical average over training samples $(x_{i},y_{i})$ of a quantity that depends on the output $g(\theta,x_{i})$ of a neural network with weights and biases collected in the parameter vector $\theta$. The function $g(\theta,x)$ can often be locally reparameterized via an architecture-dependent lifting $\phi(\theta)$, leading to the same factorization for the global loss $\ell$. This is the starting point of our analysis, captured via the following assumption:
Assumption 2.1 (Local reparameterization).
There exists $d$ and $\phi\in\mathcal{C}^{2}(\mathbb{R}^{D},\mathbb{R}^{d})$ such that: for each parameter $\theta_{0}\in\mathbb{R}^{D}$, there is an open neighborhood $\Omega\ni\theta_{0}$ and a function $f\in\mathcal{C}^{2}(\Omega,\mathbb{R})$ such that
@@
Consider $g(\theta,x)=U\sigma(V^{\top}x)$, with $\sigma(y)\coloneqq(\max(y_{i},0))_{i}$ the ReLU activation function. Denoting $\theta=(U,V)$ with $U=(u_{1},\cdots,u_{r})\in\mathbb{R}^{n\times r}$, $V=(v_{1},\cdots,v_{r})\in\mathbb{R}^{m\times r}$ (so that $D=(n+m)r$). Consider $\theta(0)=(U(0),V(0))\in\mathbb{R}^{D}$ and let $x\in\mathcal{X}_{\theta(0)}:=\mathbb{R}^{m}-\cup_{j}\{{v_{j}(0)}^{\top}x=0\}$. Then on any domain $\Omega\subset\mathbb{R}^{D}$ such that $\theta(0)\in\Omega$ and $\theta\mapsto 1(v_{j}^{\top}x>0)=\epsilon_{j,x}$ is constant over $\theta\in\Omega$, the model $g_{\theta}(x)$ can be factorized by the reparametrization $\phi_{\mathtt{ReLU}}(\theta)=(u_{j}v_{j}^{\top})_{j=1}^{r}$ (here $d=rmn$) using $g(\theta,x)=\sum_{j}\epsilon_{j,x}\phi_{j}x$, so $\ell$ can be factorized by $\phi_{\mathtt{ReLU}}$ with some forward function $f$: the reparametrization $\phi_{\mathtt{ReLU}}(\theta)$ contains $r$ matrices of size $n\times m$ (each of rank at most one, so in particular one has $d^{\prime}\leq D-r$) associated to a “local” $f$ valid in a neighborhood of $\theta$. A similar factorization is possible for deeper ReLU networks (cf Neyshabur et al. (2015); Stock & Gribonval (2022); Gonon et al. (2023)) and we still write it $\phi_{\mathtt{ReLU}}$, as further discussed in the proof of Theorem˜3.8.
@@
For an attention layer, the input $X\in\mathbb{R}^{N\times\text{dim}}$ is the horizontal concatenation of $N$ tokens $x^{(i)}\in\mathbb{R}^{\textrm{dim}}$. The layer output is
$g(\theta,X)=\mathrm{softmax}(XQ^{\top}KX^{\top})XV^{\top}O\in\mathbb{R}^{N\times\textrm{dim}}\quad\text{where}\quad\mathrm{softmax}(A)_{i}=\frac{\exp(A_{i})}{\sum_{k=1}^{N}\exp(A_{ik})},$
with $Q,K,V,O\in\mathbb{R}^{d_{1}\times\text{dim}}$. We consider the reparameterization $\phi_{\mathtt{Att}}(\theta)\coloneqq(\phi_{1},\phi_{2})$ where $\phi_{1}\coloneqq Q^{\top}K$ and $\phi_{2}\coloneqq V^{\top}O$, such that $g(\theta,X)=\mathrm{softmax}(X\phi_{1}X^{\top})X\phi_{2}$, as done in Marcotte et al. .
Thus, similarly to the linear case Example˜2.2, $L$ can be globally factorized by $\phi_{\mathtt{Att}}$ as $f$ exhibits no dependency on the specific parameter configuration $\theta_{0}$. This naturally extends to multiple attention layers by concatenating the corresponding factorizations.
@@ 2.1 Dynamics of lifted parameters: to be or not to be intrinsic?
This paper addresses a fundamental question underlying much of the efforts to characterize the implicit bias of gradient-based methods: under what conditions does the gradient flow dynamics (1) in parameter space $\theta$ lead to a dynamics on the lifted parameters $z(t)\coloneqq\phi(\theta(t))$ that can be expressed as an intrinsic gradient flow on $z$? This is notably key when attempting to establish that $z(t)$ follows a mirror flow (Gunasekar et al., 2017), which is a key step to characterize the implicit bias of gradient-based optimization. We specifically examine when $z(t)$ follows a flow with respect to a Riemannian metric which, by definition depends only on $z$ (and the initial parameter configuration $\theta_{0}$), thereby eliminating explicit dependence on the parameter trajectory $\theta(t)$.
A starting point of the analysis is that, under ˜2.1 and by the chain rule
@@
(corresponding to the so-called path kernel in when $\Phi$ is the path-lifting associated to ReLU networks Gebhart et al. (2021)) can be solely expressed in terms of $z$ and $\theta_{0}$ during the trajectory, i.e. do we have a function $K=K_{\theta_{0}}$ such that $M(\theta(t))=K(z(t))$? When this is the case (3) becomes
$\dot{z}(t)=-K(z)\nabla f(z),$ (5)
an ordinary differential equation (ODE) which is a Riemannian flow for the metric $K^{-1}(z)$ (or a sub-Riemannian flow for the pseudo-inverse $K^{+}(z)$ when $K(z)$ is not invertible) Boumal (2023), hence associated to an intrinsic dynamic on the lifted parameters $z(t)$.
As illustrated with the following example, rewriting $M(\theta(t))$ as a function of $z(t)$ along the trajectory $\theta(x)$ is indeed possible for simple linear networks, with a function $K(\cdot)$ that depends on the initialization $\theta_{0}$.
@@
Consider $g(\theta,x)=uvx$, with $\theta\coloneqq(u,v)\in\mathbb{R}_{\star}\times\mathbb{R}^{m}$, and $z=\phi(\theta)=uv\in\mathbb{R}^{m}$ (cf Example˜2.2). Then $M(\theta)=\partial\phi(\theta)\partial\phi(\theta)^{\top}=vv^{\top}+u^{2}I_{m}$. During the trajectory $u^{2}-\|v\|^{2}=u_{0}^{2}-\|v_{0}\|^{2}=:\lambda$ (as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h(\theta):=}u^{2}-\|v\|^{2}$ is conserved (Arora et al. (2019)), and as $vv^{\top}=u^{-2}zz^{\top}$ we have $(u^{2})^{2}-\lambda u^{2}-\|z\|^{2}=0$ so that $u^{2}=\frac{\lambda+\sqrt{\lambda^{2}+4\|z\|^{2}}}{2}$. As a result along the whole trajectory we have $M(\theta)=K_{\theta_{0}}(z)$ so that $z(t)$ satisfies the ODE (5) with
@@
See Appendix˜A for more comments on that example. In the above example, the function $K_{\theta_{0}}$, as its notation suggests, only depends on the initialization but not on the function $f$ such that $\ell=f\circ\phi$. In machine learning scenarios, $f$ typically captures dependence on the training dataset. The intrinsic metric $K_{\theta_{0}}(z)$ thus captures parts of the dynamics of $z(t)$ due to the network architecture (via $\phi$) and of the training algorithm (the gradient flow (1)) irrespective of the dataset and the learning task (of course the latter still play a role via the $\nabla f(z)$ term in the ODE $\dot{z}=-K_{\theta_{0}}(z)\nabla f(z)$). This motivates the introduction of the following definition.
@@
An initialisation $\theta_{0}$ verifies the intrinsic dynamic property on $\Omega$ with respect to $\phi$, if there is $K_{\theta_{0}}:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ such that, for any $f\in\mathcal{C}^{2}$, the maximal solution $\theta(\cdot)$ of (1) with $\ell=f\circ\phi$ satisfies $M(\theta(t))=K_{\theta_{0}}(\phi(\theta(t)))$ for each $t$ such that $\theta(t)\in\Omega$.

### v2

@@ Abstract
A fundamental challenge in the theory of deep learning is to understand whether gradient-based training can promote parameters belonging to certain lower-dimensional structures (e.g., sparse or low-rank sets), leading to so-called implicit bias. As a stepping stone, motivated by the proof structure of existing implicit bias analyses, we study when a gradient flow on a parameter $\theta$ implies an intrinsic gradient flow on a “lifted” variable $z=\phi(\theta)$, for an architecture-related function $\phi$. We express a so-called intrinsic dynamic property and show how it is related to the study of conservation laws associated with the factorization $\phi$. This leads to a simple criterion based on the inclusion of kernels of linear maps, which yields a necessary condition for this property to hold. We then apply our theory to general ReLU networks of arbitrary depth and show that, for a dense set of initializations, it is possible to rewrite the flow as an intrinsic dynamic in a lower dimension that depends only on $z$ and the initialization, when $\phi$ is the so-called path-lifting. In the case of linear networks with $\phi$, the product of weight matrices, the intrinsic dynamic is known to hold under so-called balanced initializations; we generalize this to a broader class of relaxed balanced initializations, showing that, in certain configurations, these are the only initializations that ensure the intrinsic metric property. Finally, for the linear neural ODE associated with the limit of infinitely deep linear networks, with relaxed balanced initialization, we make explicit the corresponding intrinsic dynamics.
@@ 1 Introduction
A central question in deep learning theory is how the complexity of gradient-based training can give rise to simpler, lower-dimensional dynamics. In this work, we explore when the gradient flow on parameters $\theta$ naturally induces a gradient flow on a “lifted” variable $z=\phi(\theta)$, where $\phi$ captures structural aspects of the model.
Intrinsic lifted flow. The study of optimization flows arising in the training of neural networks often benefits from the identification of lower-dimensional intrinsic dynamics. Specifically, due to the natural symmetries of linear and ReLU networks, it is of considerable interest to rewrite a parameter flow $\theta(t)$ in terms of an representation $z(t)=\phi(\theta(t))$, using a suitable architecture-related reparametrization $\phi$ (often called a lifting) that factors out certain symmetries.
@@
When dissected, the most advanced recent results characterizing the implicit bias induced by gradient-based optimization algorithms notably rely on two key analysis ingredients: (i) establishing that the dynamics of $z(t)$ is intrinsic, i.e., that it can be expressed as a Riemannian gradient flow with a metric depending only on $z$ and the initial parameters $\theta(0)$; (ii) further proving that this flow on $z(t)$ admits a mirror flow representation. With the combination of these two ingredients one gains access to powerful analytical tools rooted in convex optimization theory, allowing explicit characterization of the induced implicit bias. In particular, prior research has successfully leveraged mirror flow formulations to rigorously demonstrate implicit regularization effects, such as sparsity in scalar linear neural networks and two-layer networks with a single neuron (Gunasekar et al., 2018), as well as maximum-margin classification for logistic regression problems in separable data scenarios (Soudry et al., 2018; Chizat and Bach, 2020).
Recent work by Li et al. (2022) identifies sufficient conditions under which (i)+(ii) can both be established, requiring that the parametrization $\phi$ be commuting. However, this commuting condition is rarely satisfied in practical scenarios. This work focuses on characterizing weaker conditions ensuring that the flow on $z(t)$ is still driven by an intrinsic Riemannian gradient flow (but not necessarily a mirror flow anymore), which we believe is an important step forward and a starting point for future investigations encompassing variants of (ii) with warped mirror flows (Azulay et al., 2021) for practical (deep) network architectures. A first sufficient condition for (i), introduced by Marcotte et al. (2023), demands merely that the parametrization be involutive. Marcotte et al. (2023) have shown that this weaker condition applies specifically to the parametrization used in two-layer ReLU networks (Stock and Gribonval, 2022). As we will see, a consequence of the analysis conducted in our paper is the extension of this result to arbitrary DAG ReLU networks (Gonon et al., 2024).
@@
Conservation laws. The functions conserved during the training dynamics play a crucial role in establishing that the dynamics of $z(t)$ is governed by an (intrinsic) Riemannian metric that depends only on $z$ and the initialization $\theta_{0}:=\theta(0)$. Indeed, when a trajectory $\theta(t)$ is known to remain within level sets $\{\theta:\mathbf{h}(\theta)=\mathbf{h}(\theta_{0})\}$ where $\mathbf{h}$ is a (vector-valued) conserved function, the dynamics are effectively restricted to a manifold of lower dimension that is entirely determined by the initialization. A particularly important class of conserved functions along these trajectories is given by the conservation laws associated with a certain architecture-dependent parametrization $\phi$, a concept introduced in Marcotte et al. (2023). These laws depend exclusively on $\phi$, and notably, in the context of neural network training dynamics, they represent quantities preserved across trajectories irrespective of the initial conditions or the training dataset. In the specific case of linear and ReLU neural networks, these conservation laws correspond exactly to previously known “canonical” conserved functions identified in Du et al. (2018); Arora et al. (2019), as demonstrated by Marcotte et al. (2023); see also Le and Jegelka (2022) for canonical conservation laws for general DAG ReLU architectures. Furthermore, Marcotte et al. (2023) establish that if the parametrization $\phi$ is involutive, there exist sufficiently many scalar conservation laws to fully rewrite the original trajectory $\theta(t)$ in terms of $\phi(\theta(t))$ and the initial conditions alone. In the linear network case, when so-called balanced conditions (a notion introduced in Arora et al. (2019)) are satisfied (i.e., when the initialization sets all canonical conservation laws (Chitour et al., 2018) to zero, $\mathbf{h}(\theta_{0})=0$), it becomes possible to rewrite the flow in terms of $z=\phi(\theta)$, where $\phi$ corresponds to the product of weight matrices, as an intrinsic Riemannian metric (Arora et al., 2018; Bah et al., 2022). Moreover, Achour et al. (2025) extended this result to linear convolutional networks in the case of a mean squared loss, but this time for arbitrary initializations, with the Riemannian metric depending on the initialization. For linear networks and in the particular case when the loss function is the square loss, Bah et al. (2022) show that the trajectory evolves on the manifold of matrices having some fixed rank under balanced condition. Still in the square-loss setting, and in the case of two-layer linear networks, Tarmoun et al. (2021); Braun et al. (2022); Dominé et al. (2025) exploit the conservation laws to obtain an exact closed-form expression for $z(t)$ under specific configurations, whereas Varre et al. (2023) use the same laws to analyse an implicit bias of this dynamic.
@@
Our main contributions. We first define the notion of intrinsic dynamic property (Definition˜2.6), then the notion of intrinsic metric property (Definition˜2.10) and finally the one of intrinsic recoverability property (Definition˜2.15), and we show the implications (Lemma˜2.11 and Lemma˜2.16):
@@
We then provide a simple criterion that characterizes the intrinsic recoverability property (Theorem˜2.17), and show (Proposition˜2.21) that this criterion is quasi equivalent to the Frobenius property, which is slightly weaker than involutivity (Definition˜2.20). We prove that the so-called path-lifting (Gonon et al., 2024) reparametrization for general ReLU networks of arbitrary depth satisfies this property (Theorem˜3.1), characterizing a dense set of initializations of a general ReLU network that satisfies the intrinsic recoverability property (Corollary˜3.3), as illustrated by a characterization of the intrinsic dynamic of a 3-layer neural network (Proposition˜3.5). Next, by establishing a necessary condition for the intrinsic metric property to hold based on the study of kernels of linear mappings Theorem˜2.14, we show that the intrinsic metric property fails to hold for the natural reparametrizations corresponding to 2-layer linear networks (resp. of attention layers), unless the initialization satisfies the relaxed balanced condition introduced in Definition˜3.6 (Theorem˜3.9). We then show that relaxed balanced initializations do satisfy the intrinsic metric property, not only in 2-layer networks (Theorem˜3.8) but also in linear networks of arbitrary depth (Theorem˜3.11), and we characterize the resulting intrinsic dynamic. Finally, we extend our analysis to the infinite-depth limit of linear networks. We show that a set of functions is conserved along the trajectory (Proposition˜3.12), and, in contrast to the case of $L>2$-layer, we derive a closed-form expression for the metric in the case of relaxed balanced initializations (Theorem˜3.13).
@@ 2 Dynamics of over-parameterized models
In most machine learning models, overparameterization occurs due to inherent symmetries (such as rescaling) within the parameter space $\theta\in\mathbb{R}^{D}$. In practice, this redundancy can be factored out through a function $\phi$ (often called a lifting (Candès et al., 2013; Gonon et al., 2024)) that captures these symmetries. Although the resulting lifted variable $z=\phi(\theta)\in\mathbb{R}^{d}$ often lives in higher dimension $d\gg D$, it also belongs to a lower dimensional manifold $\mathcal{Z}$ of dimension $d^{\prime}<D$, and provides a representation of the essential structure of the model. We consider parameters $\theta(t)\in\mathbb{R}^{D}$ that satisfy the gradient flow dynamic with some initialization $\theta_{0}$:
$\dot{\theta}(t)=-\nabla\ell(\theta(t)),\quad\theta(0)=\theta_{0}$ (1)
to minimize the function $\ell$. In machine learning, $\ell(\theta)$ is typically defined as the empirical average over training samples $(x_{i},y_{i})$ of a quantity that depends on the output $g(\theta,x_{i})$ of a neural network with weights and biases collected in the parameter vector $\theta$. The function $g(\theta,x)$ can be locally reparameterized via the above architecture-dependent lifting $\phi(\theta)$, leading to the same factorization for the global loss $\ell$. This is the starting point of our analysis, captured via the following assumption:
Assumption 2.1 (Local reparameterization of $\ell$ via $\phi\in\mathcal{C}^{2}(\mathbb{R}^{D},\mathbb{R}^{d})$ on an open set $\Omega\subseteq\mathbb{R}^{D}$).
There is a function $f\in\mathcal{C}^{2}(\Omega,\mathbb{R})$ such that
@@
Consider $g(\theta,x)=U\sigma(V^{\top}x)$, with $\sigma(y)\coloneqq(\max(y_{i},0))_{i}$ the ReLU activation function. Denote $\theta=(U,V)$ with $U=(u_{1},\cdots,u_{r})\in\mathbb{R}^{n\times r}$, $V=(v_{1},\cdots,v_{r})\in\mathbb{R}^{m\times r}$ (so that $D=(n+m)r$). Given training vectors $x_{i}\in\mathbb{R}^{m}$ consider $\Omega\subseteq\mathbb{R}^{D}$ the set of all parameters $\theta=(U,V)$ such that $v_{j}^{\top}x_{i}\neq 0$ for every $i,j$. Then, $\theta\mapsto 1(v_{j}^{\top}x>0)=\epsilon_{j,x}$ is locally constant over $\theta\in\Omega$, and the model $g_{\theta}(x)$ can be factorized by the reparametrization $\phi_{\mathtt{ReLU}}(\theta)=(u_{j}v_{j}^{\top})_{j=1}^{r}$ (here $d=rmn$) using $g(\theta,x)=\sum_{j}\epsilon_{j,x}\phi_{j}x$ with $\phi_{j}(\theta)\coloneqq u_{j}v_{j}^{\top}$, so $\ell$ can be factorized by $\phi_{\mathtt{ReLU}}$ with some forward function $f$: the reparametrization $\phi_{\mathtt{ReLU}}(\theta)$ contains $r$ matrices of size $n\times m$ (each of rank at most one, so in particular one has $d^{\prime}\leq D-r$) associated to a “local” $f$ valid in a neighborhood of $\theta$. A similar factorization is possible for deeper ReLU networks (cf Neyshabur et al. (2015); Stock and Gribonval (2022); Gonon et al. (2024)) and we still write it $\phi_{\mathtt{ReLU}}$, as further discussed in the proof of Theorem˜3.1.
@@
For an attention layer, the input $X\in\mathbb{R}^{N\times\text{dim}}$ is the vertical stacking of $N$ tokens $x^{(i)}\in\mathbb{R}^{\textrm{dim}}$. The layer output is
$g(\theta,X)=\mathrm{softmax}(XQ^{\top}KX^{\top})XV^{\top}O\in\mathbb{R}^{N\times\textrm{dim}}\quad\text{with}\quad\bigl(\mathrm{softmax}(A)\bigr)_{ij}\coloneqq\tfrac{\exp(A_{ij})}{\sum_{k=1}^{N}\exp(A_{ik})},$
with $Q,K,V,O\in\mathbb{R}^{d_{1}\times\text{dim}}$. We use the reparameterization $\phi_{\mathtt{Att}}(\theta)\coloneqq(\phi_{1},\phi_{2})$ where $\phi_{1}\coloneqq Q^{\top}K$ and $\phi_{2}\coloneqq V^{\top}O$, such that $g(\theta,X)=\mathrm{softmax}(X\phi_{1}X^{\top})X\phi_{2}$, as done by Marcotte et al. (2025).
Thus, similarly to the linear case Example˜2.2, $\ell$ can be globally factorized by $\phi_{\mathtt{Att}}$ as $f$ exhibits no dependency on the specific parameter configuration $\theta_{0}$. This naturally extends to multiple attention layers by concatenating the corresponding factorizations.
@@ 2.1 Dynamics of lifted parameters: to be or not to be intrinsic?
This paper addresses a fundamental question underlying much of the efforts to characterize the implicit bias of gradient-based methods: under what conditions does the gradient flow dynamics equation 1 in parameter space $\theta$ lead to a dynamics on the lifted parameters $z(t)\coloneqq\phi(\theta(t))$ that can be expressed as an intrinsic gradient flow on $z$? This is crucial when attempting to establish that $z(t)$ follows a mirror flow (Gunasekar et al., 2017), which is a key step to characterize the implicit bias of gradient-based optimization. We specifically examine when $z(t)$ follows a flow with respect to a Riemannian metric which, by definition depends only on $z$ (and the initial parameter configuration $\theta_{0}$), thereby eliminating explicit dependence on the parameter trajectory $\theta(t)$.
A starting point of the analysis is that, under ˜2.1 and by the chain rule when $\theta(t)\in\Omega$
@@
(corresponding to the so-called path kernel (Gebhart et al., 2021) when $\phi$ is the path-lifting associated to ReLU networks) can be solely expressed in terms of $z$ and $\theta_{0}$ during the trajectory, i.e. do we have a function $K=K_{\theta_{0}}$ such that $M(\theta(t))=K(z(t))$? When this is the case equation 3 becomes
$\dot{z}(t)=-K(z(t))\nabla f(z(t)),$ (5)
an ordinary differential equation (ODE) which is a Riemannian flow (Boumal, 2023) for the metric $K^{-1}(z)$ (or a sub-Riemannian flow for the pseudo-inverse $K^{+}(z)$ when $K(z)$ is not invertible) , hence associated to an intrinsic dynamic on the lifted parameters $z(t)$.
As illustrated next, rewriting $M(\theta(t))$ as a function of $z(t)$ along the trajectory $\theta(t)$ is indeed possible for simple linear networks, with a function $K(\cdot)$ that depends on the initialization $\theta_{0}$.
@@
Consider $g(\theta,x)=uvx$, with $\theta\coloneqq(u,v)\in\mathbb{R}_{*}\times\mathbb{R}^{m}$, and $z=\phi(\theta)=uv\in\mathbb{R}^{m}$ (cf Example˜2.2). Then $M(\theta)=\partial\phi(\theta)\partial\phi(\theta)^{\top}=vv^{\top}+u^{2}\mathrm{Id}_{m}$. During the trajectory $u^{2}-\|v\|^{2}=u_{0}^{2}-\|v_{0}\|^{2}=:\lambda$ (as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h(\theta):=}u^{2}-\|v\|^{2}$ is conserved (Arora et al., 2019), and as $vv^{\top}=u^{-2}zz^{\top}$ we have $(u^{2})^{2}-\lambda u^{2}-\|z\|^{2}=0$ so that $u^{2}=\frac{\lambda+\sqrt{\lambda^{2}+4\|z\|^{2}}}{2}$. As a result along the whole trajectory we have $M(\theta)=K_{\theta_{0}}(z)$ so that $z(t)$ satisfies the ODE equation 5 with
@@
See Appendix˜B for more comments on that example. In the above example the function $K_{\theta_{0}}$, as its notation suggests, only depends on the initialization $\theta_{0}$ but not on the function $f$ such that $\ell=f\circ\phi$. In machine learning scenarios, $f$ typically captures dependence on the training dataset. The intrinsic metric $K_{\theta_{0}}(z)$ thus captures parts of the dynamics of $z(t)$ due to the network architecture (via $\phi$) and of the training algorithm (the gradient flow equation 1) irrespective of the dataset and the learning task (of course the latter still play a role via the $\nabla f(z)$ term in the ODE $\dot{z}=-K_{\theta_{0}}(z)\nabla f(z)$). This motivates the introduction of the following definition.
@@
$\theta_{0}$ verifies the intrinsic dynamic property with respect to $\phi$ on some open set $\Omega\ni\theta_{0}$, if there is $K_{\theta_{0}}:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ such that, for every $f\in\mathcal{C}^{2}$, the maximal solution $\theta(\cdot)$ of equation 1 on $\Omega$ with $\ell=f\circ\phi$ satisfies $M(\theta(t))=K_{\theta_{0}}(\phi(\theta(t)))$ for each $t$.